OpenStudy (daniellelovee):

Divide (x^3+2x-5)÷(x-3)

OpenStudy (anonymous):

(x^3+3x+11)+28/x-3

OpenStudy (daniellelovee):

how did you get there?

OpenStudy (anonymous):

magic

OpenStudy (daniellelovee):

-.-' I accept the joke but lets be real

OpenStudy (anonymous):

and i used synthetic division

OpenStudy (alekos):

i get x^2 + 3x + 8 + 19/x-3

OpenStudy (anonymous):

how did you get that

OpenStudy (alekos):

long division of polynomials

OpenStudy (anonymous):

is that the same as synthetic division

OpenStudy (alekos):

yes

OpenStudy (anonymous):

then i have no idea what happened and why we got different answers

OpenStudy (daniellelovee):

ok ok ppl don't argue go step by step

OpenStudy (daniellelovee):

and prguan I guess your magic just doesn't work

OpenStudy (anonymous):

alekos did the same method so "magic" does work

OpenStudy (daniellelovee):

seriously though help me solve this you two because you know karma

OpenStudy (daniellelovee):

:)

OpenStudy (alekos):

sorry. the correct answer is x^2 +3x + 11 + 28/x-3 I'll show you how its done

OpenStudy (anonymous):

so i was right the first time

OpenStudy (daniellelovee):

THANK YOU! @alekos

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

you said it not me

OpenStudy (daniellelovee):

yeah yeah so step by step @alekos? :)

OpenStudy (alekos):

x^2 _____________________ x-3 )x^3 + 0x^2 + 2x -5 x^3 - 3x^2

OpenStudy (alekos):

are you familiar with the long division of polynomials?

OpenStudy (daniellelovee):

thank you very much!

OpenStudy (alekos):

by the way prguan you were very close but you had x^3 instead of x^2 as the first term

OpenStudy (daniellelovee):

nope I'm starting on the lesson so we have to solve the "preview questions" ourselves

OpenStudy (anonymous):

3 + 2x - 5 x - 3 Determine our root divisor: To find our root, we solve the divisor equation x - 3 = 0 We add 3 to each side of the equation to get x - 3 + 3 = 0 + 3 Therefore, our root becomes x = 3 Step 1: Write down our coefficients horizontally and our root of 3 to the left: 1 0 2 -5 3 Step 2: Bring down the first coefficient of 1 1 0 2 -5 3 ↓ 1 Step 3: Multiply our root of 3 by our last result of 1 to get 3 and put that in column 2: 1 0 2 -5 3 3 1 Step 4: Add the new entry of 3 to our coefficient of 0 to get 3 and put this in the answer column 2: 1 0 2 -5 3 3 1 3 Step 5: Multiply our root of 3 by our last result of 3 to get 9 and put that in column 3: 1 0 2 -5 3 3 9 1 3 Step 6: Add the new entry of 9 to our coefficient of 2 to get 11 and put this in the answer column 3: 1 0 2 -5 3 3 9 1 3 11 Step 7: Multiply our root of 3 by our last result of 11 to get 33 and put that in column 4: 1 0 2 -5 3 3 9 33 1 3 11 Step 8: Add the new entry of 33 to our coefficient of -5 to get 28 and put this in the answer column 4: 1 0 2 -5 3 3 9 33 1 3 11 28 Our synthetic division is complete. The values in our results row form a new equation, which has a degree 1 less than our original equation shown below: Leading Answer Term = x(3 - 1) = x2 Since the last number in our result line of 28 does not equal 0, our answer will have a remainder. We take this term and divide it by our (x - r) root which is x - 3 shown below in our answer: Answer = x2 + 3x + 11 +28 __ x - 3

OpenStudy (anonymous):

HE HAD THE RIGHT ANSWER TOO

OpenStudy (anonymous):

i screwed up the first time

OpenStudy (anonymous):

now i just proved to you that "magic" works :)

OpenStudy (daniellelovee):

nvm chill no need to caplock